Multiple-Concept Example 9 reviews the concepts that are important in this problem. An electron, traveling at a speed of 5.76 × 107 m/s, strikes the target of an X-ray tube. Upon impact, the electron decelerates to one-quarter of its original speed, emitting an X-ray in the process. What is the wavelength of the X-ray photon?
Answer #1

Ei = 1/2mv2

Ef = 1/2m * (1/4v)2 = 1/32mv2

change in energy = 15/32mv2

so,

15/32mv2 = hc / \lambda

1.415e-15 = 6.63e-34 * 3e8 / \lambda

\lambda = 1.405e-10 m